example:
how would you solve this:
7740.70(1+(r/365))^365=7770.84
I can't figure out what to do with the exponent?

Answer by mukhopadhyay(490)

You will need a scientific calculator to do the necessary computation.
The steps are as below:
7740.70(1+(r/365))^365=7770.84
=> [1+(r/365)]^365 = 7770.84/7740.70
Take log on both sides:
log [1+(r/365)]^365 = log (7770.84/7740.70)
=> 365 log [1+(r/365)] = log (7770.84) - log (7740.70)
=> log [1+(r/365)] = [log (7770.84) - log (7740.70)]/365
Use the definition of the log: if log x = a, then x = 10^a
Thus, [1+(r/365)] = 10^[{log (7770.84) - log (7740.70)}/365]
=> r/365 = 10^[(log 7770.84 - log 7740.70)/(365) - 1]
=> r = 365*10^[(log 7770.84 - log 7740.70)/(365) - 1]
The above gives the exact solution of r. If you want to find the approximate solution, you will need a calculator to do so. The solution can also be found easily by using graphing calculator.
The above kind of expression looks kind of Cryptic, but you will get habituated after practicing them for a while